Saki is ice-skating to the left with a velocity of 5 m/s. A steady breeze starts blowing, causing Saki to accelerate leftward at a constant rate of 1.5m/s.

After 6s what is Saki's velocity?
Answer using a coordinate system where rightward is positive.
Answer #1

Given :

Initial velocity = u = -5 m/s

acceleration =a = -1.5 m/s2

To find :

Velocity after , t = 6s

Solution:

Here we can use

v = u + at...................(1)

where v = velocity after time t .

putting values in equation (1)

v = -5 - 1.5 \times 6 = -14 m/s.................Answer

So Saki's velocity will be 14 m/s leftwards after 6 s.